OpenStudy (anonymous):

Write the polar equation in rectangular form. (t = theta) r = 3 cos t A. x2 + y2 = 3 B. x2 + y2 = 3r C. x2 + y2 = 9 D. x2 + y2 = 3x

OpenStudy (anonymous):

you want to substitute x =rcost and y=rsint somehow into the equation if convenient. one way is multiply each side by r so you get r^2=3rcost. also remember r^2 =x^2+y^2 so simple substitution gives x^2 +y^2 =3x

OpenStudy (anonymous):

this is greatr thanks!!!!